OpenStudy (kaylak12345):

write an equation of the line that goes through the point (4,-5) and is parallelto line y=2x+8

OpenStudy (kaylak12345):

@3mar

OpenStudy (kaylak12345):

please help

OpenStudy (benji):

So we are going to use slope intercept form. So we must use the slope of the given equation and put it in a new one \[y=2x+b\] Next we need to plug the given points into our equation, we get \[-5=2(4)+b\] Solve for b and tell me what you get.

OpenStudy (3mar):

Thank you for you confidence!

OpenStudy (kaylak12345):

the slope is 2 and my parallel is y=2x-13

OpenStudy (3mar):

want me to check or answer in details?

OpenStudy (benji):

yes you were right y=2x-13 is the answer

OpenStudy (kaylak12345):

y=2x+8 slope is 2 y = mx + b slope(m) = 2 (4,-5) x = 4 and y = -5 now we sub and find b, the y intercept -5 = 2(4) + b -5 = 8 + b -5-8 = b -13 = b so parallel line/equation is : y = 2x - 13

OpenStudy (benji):

correct

OpenStudy (3mar):

Anyway, the form of the equation of a line passes through a point\[(x_1,y_1)\] and being parallel to another line is \[\frac{ y=y_1 }{ x-x_1 }=m\] where m is the slope of the given line equation.

OpenStudy (kaylak12345):

one more question

OpenStudy (3mar):

do maths, your equation is correct!

OpenStudy (kaylak12345):

3,4 -1,-4 -3,2 is it a right triangle i cant tell

OpenStudy (kaylak12345):

i think it is

OpenStudy (3mar):

Let me try

OpenStudy (kaylak12345):

ok

OpenStudy (3mar):

Take that if you want. Decide yourself. Are these points could make a right triangle?

OpenStudy (kaylak12345):

its supposed to be 3,4 lol

OpenStudy (3mar):

This is how it looks.

OpenStudy (kaylak12345):

its 3,4 lol

OpenStudy (3mar):

Sorry. modification!

OpenStudy (kaylak12345):

rhen yes lol

OpenStudy (kaylak12345):

how would i explain that it is one

OpenStudy (3mar):

If you want to check it mathematically, simply find the slopes of every segment, if you find two of them which their product equals to -1, there will be a right triangle. \[m_1*m_2=-1\] this is the condition of normal lines.

OpenStudy (kaylak12345):

so how to do i find the slope of the segments

OpenStudy (3mar):

Simply \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] Let's try the segment AB together

OpenStudy (3mar):

Hey??????????

OpenStudy (kaylak12345):

you get 2

OpenStudy (3mar):

No, brother! It is 3. Let's do it :] \[m=\frac{ 3-(-3) }{ 4-2 }=\frac{ 6 }{ 2 }=3\]

OpenStudy (kaylak12345):

but doesnyt it have to equal -1

OpenStudy (kaylak12345):

so its not a right triangle

OpenStudy (3mar):

Take the second leg of the right angle (or if it would be): slope of BC: \[m_{BC}=\frac{ -4-2 }{ -1-(-3) }=\frac{ -6 }{ 2 }=-3\]

OpenStudy (kaylak12345):

so it is a right triangle

OpenStudy (kaylak12345):

because all slopes equal the smae

OpenStudy (kaylak12345):

am i rigth

OpenStudy (kaylak12345):

if you dont hurry im goign with it

OpenStudy (3mar):

It is. How could you let me made this mistake? The slope is y/x not x/y \[m=\frac{ \Delta y }{ \Delta x } \] not \[m=\frac{ \Delta x }{ \Delta y }\]

OpenStudy (3mar):

Sorry Sorry Sorry

OpenStudy (kaylak12345):

so is it or is it not a right triangle

OpenStudy (kaylak12345):

i have to go mow soon

OpenStudy (3mar):

Two slope are 1/3 and -3. So they are perpendicular to each other. There is a right angle in the triangle. Done!

OpenStudy (kaylak12345):

so its not a right triangle